OpenStudy (onepieceftw):

Simplify (1 − sin x)(1 + sin x).

OpenStudy (mrnood):

can you write a simple form of (a+b)(a-b) use this to write your equation in a way you might recognise...

OpenStudy (onepieceftw):

Is it like square of a sum? (a+b)(a-b)=(a-b)^2

OpenStudy (er.mohd.amir):

(a-b)(a+b)=a*a-b*b= a^2-b^2

OpenStudy (mrnood):

so in your case a=1 b = sin x so rewrite the equation a^2-b^2 with those values